# Asymptotic freedom, lost: Complex conformal field theory in the two-dimensional $O(N>2)$ nonlinear sigma model and its realization in the spin-1 Heisenberg chain

### v1

@@ Abstract
and engineer the imaginary terms for $J_{2}$ and $K$ via the jump operators $\boldsymbol{L}_{i,J}=\sqrt{J_{2}^{\mathrm{Im}}}(\boldsymbol{S}_{i}-\boldsymbol{S}_{i+2})$ and $L_{i,K}=\sqrt{6K^{\mathrm{Im}}}(1-P_{0})$, where i=1,,Li=1,...,L and $P_{0}=(1/3)[(\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{i+1})^{2}-1]$ is the projector onto the singlet sector. The resulting $H_{\text{eff}}$ realizes the same Hamiltonian we studied numerically (Eq. (5)) with $J_{2}=J^{\mathrm{Re}}_{2}+iJ^{\mathrm{Im}}_{2}$ and $K=K^{\mathrm{Re}}+iK^{\mathrm{Im}}$. (There is also an additional constant imaginary shift $\delta\equiv-i(4K^{\mathrm{Im}}+J^{\mathrm{Im}}_{2})L$).
Discussion.— We have established the existence of a critical fixed point of the 2D $O(N)$ NLSM for $N>2$ at a complex value of the coupling $g$. An immediate implication is for perturbation theory: at any fixed loop order, a truncated beta function $\beta(g)$ must generically admit zeros away from the positive real axis. It would be interesting to assess whether these perturbative zeros align with the CCFT points identified here, providing a nontrivial consistency check on the perturbative expansion PhysRevLett.57.1383. More broadly, whether these complex fixed points could influence the flow on the real axis in the NLSM remains to be determined PhysRevE.107.014117.
Our microscopic analysis so far focused on the fixed point, but it would be valuable to map out the surrounding phase diagram. Based on the schematic flow in Fig. 1, one expectation is that the critical point manifests as an “unnecessary transition” within the Haldane phase, in the sense that all trajectories in its vicinity ultimately run to the same strong coupling fixed point. The situation could be richer, however, as in loop models where a CCFT describes the endpoint of a first-order line separating a “loop gas” from a “loop liquid” PhysRevLett.131.131601.
Finally, we proposed a protocol to prepare the CCFT ground state from generic initial states via continuous monitoring and postselection in a spin-1 Heisenberg chain. Our analysis so far has focused primarily on spectral properties of the corresponding non-Hermitian Hamiltonian; it would be interesting in future work to identify more directly accessible signatures of the CCFT under this dynamics, for example in correlation functions. Another especially tantalizing direction is to realize this exotic criticality directly at the master-equation level, without the need for postselection. We also note that alternative realizations of complex $O(N)$ nonlinear sigma models may be possible in monitored dynamics of free fermions PhysRevX.13.041045.
@@ Acknowledgements.
End Matter
Optimization procedure.— We use a gradient descent algorithm to locate the finite-size critical point in the parameter space $(J_{2},K)$ with the fastest finite-size convergence. We do so by matching the thermal and $\ell=1$ watermelon scaling dimensions in the spin chain spectrum with their predicted values. To extract these scaling dimensions, we compare the momentum $k$ and total spin $S$ of the eigenstates to the known symmetries of the CCFT primaries, see Refs. 10.21468/SciPostPhys.14.5.092; Gorbenko2020. We will use $E_{k,S}^{\nu}$ to denote the eigenergies and $\nu\in\mathbb{Z}^{+}$ to index the states in order of increasing real energy. The CCFT vacuum is a translationally-invariant singlet and naturally corresponds to the Haldane ground state with energy $E_{g}=E_{k=0,S=0}^{0}$. The thermal operator is a singlet, so it corresponds to $E_{k=0,S=0}^{1}$. The $\ell=1$ WM is the $O(3)$ vector operator and has momentum $k=\pi$ since this is an antiferromagnetic model: it thus corresponds to $E_{k=\pi,S=1}^{0}$. Finally, the velocity
$v=\frac{L}{2\pi}\big(E_{k=\pi+\frac{2\pi}{L},S=1}^{0}-E_{k=\pi,S=1}^{0}\big)$ (10)
can be inferred from the first descendant state of the $\ell=1$ WM at $k=\pi+2\pi/L$. The resulting $\ell=1$ WM and thermal scaling dimensions are respectively given by
$\Delta_{\ell=1}^{\text{calc}}=\frac{E_{k=\pi,S=1}^{0}-E_{g}}{2\pi v/L},\text{ and }\Delta_{\epsilon}^{\text{calc}}=\frac{E_{k=0,S=0}^{1}-E_{g}}{2\pi v/L}.$ (11)
Using the calculated scaling dimensions from exact diagonalization, we can now locate the finite-size critical point with fastest convergence by searching for the minima of the cost function $f(J_{2},K)=|R_{1}|+w|R_{2}|$ through gradient descent, where
$R_{1}=\frac{\Delta_{\ell=1}}{\Delta_{\epsilon}}-\frac{\Delta_{\ell=1}^{\text{calc}}}{\Delta_{\epsilon}^{\text{calc}}},\text{ }R_{2}=\frac{\Delta_{\epsilon}}{\Delta_{\epsilon}^{\text{calc}}}-1.$ (12)
and $w=0.2$ is a feature scaling weight which we found to help the gradient descent algorithm. We emphasize that the two-parameter optimization over $J_{2}$ and $K$ is used solely to improve finite-size numerics, and only a single complex tuning parameter should be needed to realize the CCFT in the thermodynamic limit.
Identification of watermelon operators in the spectrum.—We now discuss the identification of $\ell$-leg watermelon operators in our spectrum for $\ell=1,2,3,4$. For each $\ell$ value, we find the decomposition into $O(N)$ irreducible representations in Ref. 10.21468/SciPostPhys.14.5.092, and we then specialize to $N=3$ (for a reference on the multiplicity of Young tableaux of $O(N)$, see Cvitanovic2008). We find perfect agreement with the multiplicity we observe in our numerics (see clusters labeled A, B, C, D) in Fig. 3c). For $\ell=1$, Ref. 10.21468/SciPostPhys.14.5.092 gives the $O(N)$ irreducible representation $[1]$ (using row-length notation for Young tableaux), i.e. an $S=1$ state in the case of $N=3$. This is the $O(N)$ vector field. We confirm this numerically (point A). For $\ell=2$, Ref. 10.21468/SciPostPhys.14.5.092 gives $[2]$, i.e. an $S=2$ state, which we confirm numerically (point B). This is the symmetric traceless rank-2 tensor. For $\ell=3$, Ref. 10.21468/SciPostPhys.14.5.092 gives $[3]+[111]$, which for $N=3$ specializes to $3\oplus 0$, i.e. a spin-3 and a singlet. Microscopically, the singlet operator is identified with $\boldsymbol{S}_{i}\cdot(\boldsymbol{S}_{j}\times\boldsymbol{S}_{k})$. We confirm the presence of both the spin 0 (point C0) and the spin 3 (point C3) states. For $\ell=4$, Ref. 10.21468/SciPostPhys.14.5.092 gives $[4]+[2]+[0]+[22]+[211]$. For $O(3)$, the multiplicity of $[22]$ and $[211]$ is zero (see e.g. Table 10.3 and 10.4 of Cvitanovic2008), leaving only $4\oplus 2\oplus 0$. The singlet state is the previously discussed loop crossing operator that appears in the action and is weakly irrelevant. We have confirmed the presence of a $S=4$, $S=2$ and $S=0$ state close to the predicted $\ell=4$ WM scaling dimension (see points D0, D2, and D4). We summarize all the calculated scaling dimensions in Fig. 3(d).
Entanglement entropy and central charge.—We provide further evidence of the CCFT by calculating the central charge through two complementary approaches: the entanglement entropy and the finite-size scaling of the ground state energy.
For non-Hermitian Hamiltonians, the generalized entanglement entropy is a complex quantity n578-ljd5; PhysRevResearch.2.033069; PhysRevLett.128.010402; Guo_2021; https://doi.org/10.48550/arxiv.2403.03259; PhysRevLett.119.040601; 10.21468/SciPostPhys.12.6.194; PhysRevLett.130.241602; 10.21468/SciPostPhysCore.6.3.062; PhysRevB.107.205153. It is defined using the biorthonormal ground-state density matrix $\rho=|\psi_{R}\rangle\langle\psi_{L}|$, where $|\psi_{R}\rangle$ and $\langle\psi_{L}|$ denote the right and left ground states satisfying $H|\psi_{R}\rangle=E_{g}|\psi_{R}\rangle$ and $\langle\psi_{L}|H=\langle\psi_{L}|E_{g}$. The entanglement entropy $S(l)=\mathrm{Tr}_{A}(\rho_{A}\log\rho_{A})$ follows from the reduced density matrix $\rho_{A}=\mathrm{Tr}_{B}(\rho)$, obtained by tracing out a subsystem $B$ of length LlL-l. At criticality, $S(l)$ displays the characteristic logarithmic scaling controlled by the central charge $c$,
$S(l)=\frac{c}{3}\log\left(\frac{L}{\pi}\sin\frac{\pi l}{L}\right)+s_{0},$ (13)
where $s_{0}$ is a non-universal constant and $c$ may be complex n578-ljd5; 10.21468/SciPostPhys.12.6.194; Calabrese_2004; HOLZHEY1994443; PhysRevLett.90.227902; PhysRevLett.104.095701. The numerical results, calculated using the density matrix renormalization group (DMRG) on a length $L=14$ chain with periodic boundary conditions PhysRevLett.69.2863; 10.21468/SciPostPhysCodeb.4, are presented in Figs. 5(a)–(b) and display agreement with the logarithmic scaling form. The optimal fit on the interval $l\in[2,12]$ yields c=1.5290.161ic=1.529-0.161i within $2\%$ of the predicted value $c_{+}=1.51-0.158i$.
The central charge $c$ also governs the scaling of the ground-state energy as
$E_{g}=e_{\infty}L-\frac{\pi vc}{6L}+O(L^{-2}),$ (14)
where $e_{\infty}$ denotes the non-universal bulk energy density and $v$ is the velocity PhysRevLett.56.742; DiFrancesco1997; https://doi.org/10.48550/arxiv.0807.3472. In Fig. 5(c)–(d), we plot $E_{g}/L$ and fit the data using $e_{\infty}$ as a free parameter. Here, we have used the value of $c$ obtained from the entanglement entropy and the velocity determined from the spectrum (see Eq. (10)). The agreement indicates that the finite-size scaling is consistent with the CCFT prediction.
Figure 5: (a)–(b) Real and imaginary parts of the complex entanglement entropy $S_{l}$ for different subsystem sizes $l$ at the CCFT. The dashed line indicates the fit to the logarithmic scaling expected at criticality [see Eq. (13)], which estimates the central charge to be c=1.5290.161ic=1.529-0.161i. (c)–(d) Real and imaginary parts of the ground-state energy density $E_{g}/L$ and fit to $E_{g}/L=e_{\infty}-\pi vc/(6L^{2})$, where $e_{\infty}$ is a fitting parameter, $c$ is determined from the entanglement entropy [see panels (a)–(b)], and $v$ is obtained from the energy spectrum [see Eq. (10)].

### v2

@@ Abstract
and engineer the imaginary terms for $J_{2}$ and $K$ via the jump operators $\boldsymbol{L}_{i,J}=\sqrt{J_{2}^{\mathrm{Im}}}(\boldsymbol{S}_{i}-\boldsymbol{S}_{i+2})$ and $L_{i,K}=\sqrt{6K^{\mathrm{Im}}}(1-P_{0})$, where i=1,,Li=1,...,L and $P_{0}=(1/3)[(\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{i+1})^{2}-1]$ is the projector onto the singlet sector. The resulting $H_{\text{eff}}$ realizes the same Hamiltonian we studied numerically (Eq. (5)) with $J_{2}=J^{\mathrm{Re}}_{2}+iJ^{\mathrm{Im}}_{2}$ and $K=K^{\mathrm{Re}}+iK^{\mathrm{Im}}$. (There is also an additional constant imaginary shift $\delta\equiv-i(4K^{\mathrm{Im}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}J^{\mathrm{Im}}_{2})L$).
Discussion.— We have established the existence of a critical fixed point of the 2D $O(N)$ NLSM for $N>2$ at a complex value of the coupling $g$. An immediate implication is for perturbation theory: at any fixed loop order, a truncated beta function $\beta(g)$ must generically admit zeros away from the positive real axis. It would be interesting to assess whether these perturbative zeros align with the CCFT points identified here, providing a nontrivial consistency check on the perturbative expansion Bernreuther and Wegner (1986), see recent discussion in Ref. Yamamoto and Kawabata (2026). More broadly, whether these complex fixed points could influence the flow on the real axis in the NLSM remains to be determined Burgelman et al. (2023).
Our microscopic analysis so far focused on the fixed point, but it would be valuable to map out the surrounding phase diagram. Based on the schematic flow in Fig. 1, one expectation is that the critical point manifests as an “unnecessary transition” within the Haldane phase, in the sense that all trajectories in its vicinity ultimately run to the same strong coupling fixed point. The situation could be richer, however, as in loop models where a CCFT describes the endpoint of a first-order line separating a “loop gas” from a “loop liquid” Haldar et al. (2023).
Finally, we proposed a protocol to prepare the CCFT ground state from generic initial states via continuous monitoring and postselection in a spin-1 Heisenberg chain. Our analysis so far has focused primarily on spectral properties of the corresponding non-Hermitian Hamiltonian; it would be interesting in future work to identify more directly accessible signatures of the CCFT under this dynamics, for example in correlation functions. Another especially tantalizing direction is to realize this exotic criticality directly at the master-equation level, without the need for postselection. We also note that alternative realizations of complex $O(N)$ nonlinear sigma models may be possible in monitored dynamics of free fermions Fava et al. (2023).